OpenStudy (anonymous):

Solve. 3x2 = 33x + 24!!??

jimthompson5910 (jim_thompson5910):

3x^2 = 33x + 24 3x^2 - 33x - 24 = 0 Now use the quadratic formula to solve for x

jimthompson5910 (jim_thompson5910):

The quadratic formula is \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] and in this case a = 3 b = -33 c = -24

OpenStudy (anonymous):

hmmm i am really bad at thiss!?? :(

jimthompson5910 (jim_thompson5910):

ok I'll take it a bit further

OpenStudy (anonymous):

thank you so much :)

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-33)\pm\sqrt{(-33)^2-4(3)(-24)}}{2(3)}\] \[\Large x = \frac{33\pm\sqrt{1089-(-288)}}{6}\] \[\Large x = \frac{33\pm\sqrt{1089+288}}{6}\] Let me know if you can take it from here or not

OpenStudy (anonymous):

woww your good at this!! im still stuck though if you could just take it to the end for me i would be forever grateful :))))

jimthompson5910 (jim_thompson5910):

Alright, \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-33)\pm\sqrt{(-33)^2-4(3)(-24)}}{2(3)}\] \[\Large x = \frac{33\pm\sqrt{1089-(-288)}}{6}\] \[\Large x = \frac{33\pm\sqrt{1089+288}}{6}\] \[\Large x = \frac{33\pm\sqrt{1377}}{6}\] \[\Large x = \frac{33+\sqrt{1377}}{6} \ \text{or} \ x = \frac{33-\sqrt{1377}}{6}\] \[\Large x = \frac{33+9\sqrt{17}}{6} \ \text{or} \ x = \frac{33-9\sqrt{17}}{6}\] \[\Large x = \frac{3(11+3\sqrt{17})}{6} \ \text{or} \ x = \frac{3(11-3\sqrt{17})}{6}\] \[\Large x = \frac{11+3\sqrt{17}}{2} \ \text{or} \ x = \frac{11-3\sqrt{17}}{2} \]

jimthompson5910 (jim_thompson5910):

The last line is one way to write the fully simplified exact answers the approximate answers (after using a calculator) are \[\Large x \approx 11.684658 \ \text{or} \ x \approx -0.684658\]

OpenStudy (anonymous):

wowww thank you sooooo much!!! i am forever grateful now haha :))))))

jimthompson5910 (jim_thompson5910):

lol glad to be of help

OpenStudy (anonymous):

yess you are best math teacher on here!!!!

jimthompson5910 (jim_thompson5910):

thanks lol

OpenStudy (anonymous):

:))

OpenStudy (anonymous):

ohhh noo we have a problem jim!!

OpenStudy (anonymous):

there is no answer like that out of the choices\ :(

jimthompson5910 (jim_thompson5910):

ok if possible show me the answer choices with a screenshot

OpenStudy (anonymous):

okk :)

OpenStudy (anonymous):

ahh turns out you answered this question for somebody else so its A!

OpenStudy (anonymous):

so thanks again!

jimthompson5910 (jim_thompson5910):

oh glad you figured it out, I'm betting the answer is probably in a different format or something

OpenStudy (anonymous):

exactly! :)

jimthompson5910 (jim_thompson5910):

ah probably something like \[\Large x = \frac{11}{2}+\frac{3\sqrt{17}}{2} \ \text{or} \ x = \frac{11}{2}-\frac{3\sqrt{17}}{2} \]

OpenStudy (anonymous):

yeap :) great job!!!!

jimthompson5910 (jim_thompson5910):

hmm clever lol, hate when it gets that picky though

OpenStudy (anonymous):

mee tooo im already confused enough haha

jimthompson5910 (jim_thompson5910):

yeah it should be smart enough to know that \[\Large x = \frac{11+3\sqrt{17}}{2} \ \text{or} \ x = \frac{11-3\sqrt{17}}{2} \] is the same as \[\Large x = \frac{11}{2}+\frac{3\sqrt{17}}{2} \ \text{or} \ x = \frac{11}{2}-\frac{3\sqrt{17}}{2} \]